OpenStudy (anonymous):

1/2x+1/3y=3 Which of the following equations is equivalent to the given equation? 3x + 2y = 3 2x + 3y = 18 3x + 2y = 18

OpenStudy (anonymous):

The equation is, (1/2)x+(1/3)y=3 =>x/2+y/3=3 =>(3x+2y)/(3)(2)=3 =>3x+2y=3(3)(2)=9(2)=18 Now you find yourself.